Can someone show me how to find the equation for the line going through.
(2,3), (5,11) and (3,1) (3, -1)

Answer by stanbon(75887)

find the equation for the line going through.
(2,3), (5,11)
slope = (11-3)/(5-2) = 8/3
Form: y = mx+b
you know x,y, and m; solve for b
3 = (8/3)2 + b
3 = 16/3 + b
b = -7/3
-------------
y = (8/3)x -(7/3
====================
and (3,1) (3, -1)
slope = (-1-1)/(3-3) = -2/0 which is undefined
The equation is x=3
=====================
Cheers,
Stan H.